find the total areaof a solid cubeof volume 64cm^3
Answer by stanbon(75887)

find the total area of a solid cubeof volume 64cm^3
----
Volume = (side)^3
---
Total area = 6(side)^2
-------
Solve: s^3 = 64 cm^3
s = 4 cm
----
Total area = 6*4^2 = 6*16 = 96 cm^2
===============================
Cheers,
Stan H.
=============